Problem 10.65

In an amusement park water slide, people slide down an essentially frictionless tube. The top of the slide is 3.2m above the bottom where they exit the slide, moving horizontally, 1.2m above a swimming pool.

Part A

What horizontal distance do they travel from the exit point before hitting the water?

Express your answer to two significant figures and include the appropriate units.

d =
Answer #1

mgh = 1/2 *mv^2

v = sqrt (2gh

= 7.92 m/s

**************

s = 1/2*a*t^2

1.2 = 0.5 *9.8 *t^2

t = 0.49

d = v*t

distance = 3.92 m